OpenStudy (anonymous):

Solve by factoring. 4x2 - 5x + 1 = 0

OpenStudy (anonymous):

\[4x^2-5x+1=(x-1)(4x-1)\]

OpenStudy (anonymous):

split it up into 4x^2 - 4x - x + 1

OpenStudy (anonymous):

so to solve you have \[(x-1)(4x-1)=0\] \[x-1=0\] \[x=1\] or \[4x-1=0\] \[4x=1\] \[x=\frac{1}{4}\]